OpenStudy (anonymous):

Solve |2x + 1 / 3| = 4 X= 2 or X = 5 X = 11/2 or -13 / 2 X = -11/2 or -13/2 X= -11/2 or 13/2

OpenStudy (anonymous):

is this \[|2x+\frac{1}{3}|==4\] or \[|\frac{2x+1}{3}|=4\]?

OpenStudy (anonymous):

The second one @satellite73

OpenStudy (anonymous):

multiply both sides by 3 first to get rid of the fraction leave the absolute value sign there what do you get?

OpenStudy (anonymous):

@sateillte73 I got 7 and 12

OpenStudy (anonymous):

i get \[|2x+1|=12\]

OpenStudy (anonymous):

then solve two equations \[2x+1=12\] and \[2x+1=-12\]

OpenStudy (anonymous):

@sateillte73 I got 11 and -13

OpenStudy (anonymous):

you get \[2x=11\] or \[2x=-13\] you still have to solve for \(x\)

OpenStudy (anonymous):

What would I plug in for X ? Can I even solve for x ?

OpenStudy (anonymous):

@sateillte73

OpenStudy (kdr1308):

You divide both sides by 2 for each one

OpenStudy (anonymous):

Okay that's what I thought @Kdr1308

OpenStudy (anonymous):

So it would be 11/2 or -13/2 or -11/2 or -13/2 ?

OpenStudy (anonymous):

@Kdr1308

OpenStudy (anonymous):

Okay that's what I thought @Kdr1308